Laser temperature sensors fixed

Go down

Laser temperature sensors fixed

Post  Alessio on Wed Nov 11, 2009 3:00 pm

Yesterday (with F. Paoletti), during the maintenance, we went to the TCS benches to try to solve the issue of the lasers temperature sensors: the output signal not being stable when touching related cables and/or connectors.
We found two different problems and fixed both.

First problem: between the temperature cable from the laser body and the electronics, there was an extension cable. The connection between the temperature cable and the extension was very loose and the temperature was changing by several degrees by touching this connection. We removed the extension cable by changing the pin layout on the electronics. The connection is now robust and there is no temperature change when touching it.

Second problem: the temperature cable is a very good RF antenna. So whenever touching the electronics and/or the connector, it is possible to change the coupling to the RF signal, thus changing the temperature reading. This problem also was solved by putting one ferrite on each temperature cable of each laser.
The result is that now we can be confident that we are reading the real temperature of the lasers. The temperatures of the lasers are about 18.5°C and are more compatible with the temperatures of the chillers that are about 14°C.

These operations required, for safety reasons, to switch off the TCS lasers and chillers one at a time. When we turned on again the NI chiller, a problem of communication occured with the serial/ethernet switch. It was solved by F. Carbognani by rebooting the device.
avatar
Alessio

Posts : 127
Join date : 2008-11-07
Age : 46
Location : Rome

View user profile

Back to top Go down

Back to top

- Similar topics

 
Permissions in this forum:
You cannot reply to topics in this forum